OpenStudy (df001):

For each value of Θ, use half angle formula find Sin(1/2)Θ Θ=300 degrees

OpenStudy (df001):

please help me!!! I need to know this now

OpenStudy (anonymous):

cos(theta)=1-2sin(theta/2)^2 therefore applying theta=300 we get cos(theta)=1/2 so, 1/2=1-2sin(theta/2)^2 evaluating this we will get, sin(theta/2)=1/2